penny spent 3/8 of her monthly salary on rent, and 1/3 of her monthly salary on food. If she has $294 left, what is her monthly salary.
Answer by venugopalramana(3286)

penny spent 3/8 of her monthly salary on rent, and 1/3 of her monthly salary on food. If she has $294 left, what is her monthly salary.
LET HER MONTHLY SALARY = X $
3/8 TH OF IT =X*3/8=3X/8 $ SPENT ON RENT
1/3 RD.OF IT=X/3 $ SPENT ON FOOD
TOTAL SPENT = 3X/8 + X/3=(9X+8X)/6=17X/24
REMAINING =X-17X/24=7X/24=294
X=24*294/7=1008